OpenStudy (anonymous):

Fun problem of the day, Find the sum of the series \[ \frac 1 2 + \frac 3 4 + \frac 5 8 +\frac 7 {16} \cdots \] Enjoy!

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{2n-1}{2^n}=\sum_{n=1}^{\infty}\frac{n}{2^{n-1}}-\sum_{n=1}^{\infty}\frac{1}{2^{n}}\]

OpenStudy (anonymous):

For the second part,\[\sum_{n=0}^{\infty}ar^n=\frac{a}{1-r}\]Note the 0 on the bottom, take away the term for n=0 for your example.

OpenStudy (anonymous):

Oh, sorry, you're not asking!